OpenStudy (anonymous):

Condense the expression to the logarithm of a single quantity. 3logx+4logy-6logz

OpenStudy (anonymous):

logx^3-logy^4-logz^6 = log(x^3 . y^4) - logz^6 = log[ (x^3 . y^4)/ (z^6)]

OpenStudy (cj7529):

3logx + 4logy - 6logz =log(x^3) + log(y^4) - log(z^6) =log(x^3 y^4) - log(z^6) =log[ (x^3 y^4)/z^6 ]

OpenStudy (anonymous):

thank you thats what i got